OpenStudy (anonymous):

need help H5P1 AND H5P3

OpenStudy (gorv):

@afridi112 H5P1 VIN -0.5 RL 8000 VIN for saturation =0.093

OpenStudy (gorv):

@afridi112 H5P3 vout=(1-1/(sqrt(1+2*RS*K*(VIN-VT))))*vin vo=0.00328 vo/vi=0.939

OpenStudy (anonymous):

What formula did you use to calculate vo? I tried with vo = -1*(VIN-VT)*vi *(Rs*K) but doesn't work.